# Charged analogues of a singularity-free anisotropic compact star under linear $f(Q)$- action

### v3

@@ 5.3 Causality condition
In the pursuit of a realistic model for an anisotropic compact star, a valuable approach for characterising the dense interior matter lies in the investigation of sound wave velocities. The radial and tangential sound velocities are expressed as $v_{r}^{2}=\frac{dp_{r}}{d\rho}$ and $\frac{dp_{t}}{d\rho}$, respectively, where $\rho,p_{r}$ and $p_{t}$ are the energy density, and radial and tangential pressures, respectively. In this particular formulation, we have used the system of units $h=c=1$, and the causality criterion imposes an absolute upper bound of the sound velocities as $v_{r}^{2}\leq 1$ and $v_{t}^{2}\leq 1$. However, the thermodynamic equilibrium ensures that $v_{r}^{2}>0$ and $v_{t}^{2}>0$. Hence, the combined effects of both conditions, i.e., $0<v_{r}^{2}\leq 1$ and $0<v_{t}^{2}\leq 1$, must be simultaneously satisfied within the stellar configuration. To circumvent the mathematical intricacies, we have opted to represent the radial variation of sound velocities graphically in Figs. 9 and 10.
Figure 9: Radial variation of radial sound velocity $(v_{r}^{2})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
Figure 10: Radial variation of tangential sound velocity $(v_{t}^{2})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
It is evident here that the causality conditions are well maintained within the stellar interior.
@@ 5.4 Energy conditions
In the field of gravitational theory, constraints known as energy conditions are applied to matter distributions. These constraints are essential for deriving a physically feasible energy-momentum tensor. Essentially, these conditions provide a method to explore the characteristics of matter distribution without requiring detailed information about the intrinsic matter content. Therefore, it is feasible to discern the physical characteristics of extreme phenomena, such as gravitational collapse or the presence of a geometrical singularity, without explicit knowledge of energy density or pressure. Essentially, the examination of energy conditions is an algebraic issue [77], more precisely, it is an eigenvalue problem associated with the energy-momentum tensor. In a 4D spacetime, the investigation of energy conditions results in the roots of a quartic polynomial, a process that becomes complex due to the existence of analytical solutions for eigenvalues. Although finding a general solution can be challenging, a physically plausible fluid distribution should simultaneously adhere to the null, weak, strong, and dominant energy conditions [77, 78, 79] within the confines of the stellar boundary. In this study, we have examined the energy conditions [80, 81] for the current stellar configuration in the following form:
@@
Figure 11: Radial variation of energy conditions for different charge intensities $(q_{0})$. Here, Red, Green, Blue, Orange, Magenta and Brown lines represent $(\rho+\frac{q^{2}}{8\pi r^{4}})$, $(\rho+p_{r})$, $(\rho+p_{t}+\frac{q^{2}}{4\pi r^{4}})$, $(\rho+p_{r}+2p_{t}+\frac{q^{2}}{4\pi r^{4}})$, $(\rho-p_{r}+\frac{q^{2}}{4\pi r^{4}})$ and $(\rho-p_{t})$ respectively. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@ 6 Stability analysis
The stability of this model is examined using the following methodologies:
@@
    Generalised TOV equation
@@
6.1 Generalised TOV equation
Examining the stability of a model subjected to various forces is essential. For an anisotropic charged compact object, this stability analysis focuses on four force components: (i) the gravitational force $(F_{g})$, (ii) the hydrostatic force $(F_{h})$, (iii) the anisotropic force $(F_{a})$, and (iv) the electromagnetic force $(F_{q})$. To ensure that the model remains in stable equilibrium, the combined effects of these forces must be balanced. In this study, to analyse the stability, we utilise the generalised form of the Tolman–Oppenheimer–Volkoff (TOV) equation [69, 70], as expressed below:
$-\frac{M_{G}(\rho+p_{r})}{r^{2}}e^{\lambda-\nu}-\frac{dp_{r}}{dr}+\frac{2\Delta}{r}+\frac{q}{4\pi r^{4}}\frac{dq}{dr}=0,$ (45)
where, $M_{G}$ is demarcated as the active gravitational mass obtained from the Tolman-Whittaker mass formula [82] in the following form:
@@ 6.1 Generalised TOV equation
$-\nu^{\prime}(\rho+p_{r})-\frac{dp_{r}}{dr}+\frac{2\Delta}{r}+\frac{2\Delta}{r}+\frac{q}{4\pi r^{4}}\frac{dq}{dr}=0.$ (47)
@@
Figure 12: Radial variation of different forces for different charge intensities $(q_{0})$. Here, Green, Red, Blue and Magenta and Brown lines represent $F_{g}$, $F_{h}$, $F_{a}$ and $F_{q}$ respectively. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@ 6.2 Herrera cracking condition
According to Herrera’s work [83], the concept of “Cracking” or “Breaking” in the context of a self-gravitating compact object refers to the appearance of total radial forces in various regions of the spherical structure, each with different signs, when perturbations are introduced into the system. If the EoS is taken into account, cracking can arise from two scenarios, namely, i) a local anisotropic fluid, or ii) a slowly contracting and radiating perfect fluid. Utilising this cracking condition, Abreu et al. [84] explored the impact of local anisotropy on the distribution of local anisotropic fluid. They demonstrated that when perturbations exist in the system, potential stability can be established through the difference between the squares of radial $(v_{r}^{2})$ and tangential $(v_{t}^{2})$ sound velocities expressed as:
$0\leq|v_{r}^{2}-v_{t}^{2}|\leq 1$ (52)
A stellar model that adheres to this relation is referred to as a stable structure. From Fig. (13), we note that the Abreu inequality is well maintained throughout the interior of the stellar configuration.
Figure 13: Radial variation of $|v_{r}^{2}-v_{t}^{2}|$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@ 6.3 Adiabatic index
The adiabatic index $(\Gamma)$ is the ratio of two specific heats that signify the rigidity of the EoS at a particular density. In the context of a relativistic anisotropic stellar structure, the adiabatic index is represented as:
@@
$\Gamma>\Gamma^{\prime}_{max},$ (54)
@@
$\Gamma^{\prime}_{max}=\frac{4}{3}-\Bigg{[}\frac{4}{3}\frac{(p_{r}-p_{t})}{|p^{\prime}_{r}|r}\Bigg{]}_{max}.$ (55)
Now, to determine the anisotropic limit, we have evaluated $\Gamma^{\prime}_{max}$ for different charge intensities $(q_{0})$ and the composite results are plotted in Fig. 14.
Figure 14: Radial variation of adiabatic index $(\Gamma)$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@ 6.4 Zel’dovich condition
Figure 15: Radial variation of $(\Omega_{r})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
Figure 16: Radial variation of adiabatic index $(\Omega_{t})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006$ respectively.
@@ 7 Discussion
Finally, we summarise the major findings of the present paper and our efforts to mimic a charged compact stellar structure in the framework of symmetric teleparallel $f(Q)$ gravity. Taking into account a spherically symmetric space-time, we combine the KB metric ansatz [62], with a linear form of the $f(Q)$ action, which allows us to formulate a singularity-free solution of the Einstein-Maxwell field equations within the framework of $f(Q)$ theory. This solution incorporates a specific form of charge, denoted as $q(r)=q_{0}r^{3}$, as outlined in Eq. (23), where $q_{0}$ is the charge intensity and $q_{0}=0$ illustrates an uncharged stellar configuration.
Within this theoretical construct, we have selected PSR J0740+6620 [73] as a potential candidate for a charged compact star. We analyse and represent our results, both numerically and graphically, by varying the charge intensity $(q_{0})$. It has been shown in Refs. [63, 64] that the modified theories of gravity lead to alterations in the boundary conditions of GR. This paper addresses these changes by assessing the continuity of both the induced metric and the extrinsic curvature tensors at the boundary of the stellar structure. We have used, i) the matching of interior solutions with the Reissner-Nordström [65, 66] exterior to compute the extrinsic curvature tensors, as obtained in Eqs. (39) and (40), and ii) the condition of nullification of radial pressure at the stellar boundary, to compute the necessary constant present in the model. We have noted the following key characteristic features:
@@
    Radial variation of the thermodynamic properties of a stellar structure, such as the energy density $(\rho)$, radial $(p_{r})$ and tangential $(p_{t})$ pressure profiles are represented in Figs. 1, 2, 3. As the charge intensity increases, the energy density and pressure profiles decrease. However, the monotonically decreasing nature of the characteristic profiles is well preserved in this model. Furthermore, the radial variation of anisotropy is illustrated in Fig. 4. From Fig. 4, we note that, $\Delta>0$, which indicates a repulsive anisotropic behaviour. This nature further indicates a more massive stellar structure. We also note that at the centre $(r\rightarrow 0)$, $\Delta\rightarrow 0$, i.e., $p_{r}=p_{t}$, which is another viable condition in favour of stable compact star modeling. The radial variations of pressure and density gradients are negative, as shown in Figs. 5, 6 and 7, which aids in the maximisation of the core density. Hence, it may be concluded that the physical parameters are consistent throughout the stellar structure.
@@
    Obtaining the maximum mass through TOV equations depends on the particular choice of the EoS. In the present scenario, we have obtained the suitable EoS by maximising the radial sound velocity at the stellar centre, i.e., $\frac{dp_{r}}{d\rho}|_{r=0}=1$. Moreover, we have used the method of curve fitting to obtain the best fit EoS in this model, which are tabulated in Table 1 for increasing charge intensity $(q_{0})$. Using the EoS, we numerically evaluate the TOV equation, and Fig. 8 depicts the mass-radius plot. The maximum masses and radii for the increasing charge intensity are listed in Table 2. We show the predicted radii of different recently observed pulsars and GW 190814 in Table 3 and it is noted that our model complies with the observational results of compact star candidates.
@@
    Figs. 9 and 10 show the variations of radial $(v_{r}^{2})$ and tangential $(v_{t}^{2})$ sound velocities with the radial distance.Notably, with increasing charge intensity, both $v_{r}^{2}$ and $v_{t}^{2}$ decrease within the stellar boundary. However, it is evident that both the velocities are well-regulated, as they do not surpass the speed of light and comply with the causality condition.
@@
    In the framework of $f(Q)$ gravity and considering the presence of charge, we have studied the necessary and sufficient energy conditions for a well-behaved compact star model, which is demonstrated in Fig. (11). We note that, all the energy conditions are well maintained throughout the stellar structure.
@@ 8 Acknowledgments
DB is thankful to DST, Govt. of India, for providing the fellowship vide no: DST/INSPIRE/03/2022/000001.

### v4

@@ 5.3 Causality condition
In the pursuit of a realistic model for an anisotropic compact star, a valuable approach for characterising the dense interior matter lies in the investigation of sound wave speeds. The radial and tangential sound speeds are expressed as $v_{r}^{2}=\frac{dp_{r}}{d\rho}$ and $\frac{dp_{t}}{d\rho}$, respectively, where $\rho,p_{r}$ and $p_{t}$ are defined earlier. In this formulation, we have used $\hbar=c=1$, and the causality criterion imposes an absolute upper bound of the sound velocities as $v_{r}^{2}\leq 1$ and $v_{t}^{2}\leq 1$. However, the thermodynamic equilibrium ensures that $v_{r}^{2}>0$ and $v_{t}^{2}>0$. Hence, the combined effects of both conditions, i.e., $0<v_{r}^{2}\leq 1$ and $0<v_{t}^{2}\leq 1$, must be simultaneously satisfied within the stellar configuration. To circumvent the mathematical intricacies, we have opted to represent the radial variation of sound velocities graphically in Figs. 9 and 10.
Figure 9: Radial variation of radial sound velocity $(v_{r}^{2})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
Figure 10: Radial variation of tangential sound velocity $(v_{t}^{2})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
It is evident here that the causality conditions are well maintained in the stellar interior within the parameter space used here.
@@ 5.4 Energy conditions
In the field of gravitational theory, constraints known as energy conditions are applied to matter distributions. These constraints are essential for deriving a physically feasible energy-momentum tensor. Essentially, these conditions provide a method to explore the characteristics of distribution of matter without requiring detailed information about the composition of internal matter. Therefore, it is feasible to discern the physical characteristics of extreme phenomena, such as the presence of a geometrical singularity or gravitational collapse of merger, without explicit knowledge of pressure or energy density. Essentially, the examination of energy conditions is an algebraic issue [77], more precisely, it is an eigen value problem associated with the energy-momentum tensor. In a 4D spacetime, the investigation of energy conditions results in the roots of a quartic polynomial, a process that becomes complex due to the existence of analytical solutions for eigenvalues. Although finding a general solution can be challenging, a physically plausible fluid distribution should simultaneously adhere to the dominant energy, strong, weak and null conditions commonly known as the energy conditions [77, 78, 79] within the confines of the stellar configuration. In this study, we have examined the energy conditions [80, 81] for the current stellar configuration in the form given below:
@@
Figure 11: Radial variation of energy conditions for different charge intensities $(q_{0})$. Here, Red, Green, Blue, Orange, Magenta and Brown lines represent $(\rho+\frac{q^{2}}{8\pi r^{4}})$, $(\rho+p_{r})$, $(\rho+p_{t}+\frac{q^{2}}{4\pi r^{4}})$, $(\rho+p_{r}+2p_{t}+\frac{q^{2}}{4\pi r^{4}})$, $(\rho-p_{r}+\frac{q^{2}}{4\pi r^{4}})$ and $(\rho-p_{t})$ respectively. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@ 6 Stability analysis
TO examine the stability of stellar structure in this model, we have applied the following methodologies:
@@
    TOV equation in a generalised form in presence of charge
@@
6.1 TOV equation in a generalised form in presence of charge
A model is physically realistic or not can be verified using the stability analysis subjected to various forces associated with the stellar configuration. For a charged compact star in presence of pressure anisotropy, this stability analysis focuses on four different components of forces: (i) gravitational force $(F_{g})$, (ii) hydrostatic force $(F_{h})$, (iii) anisotropic force $(F_{a})$, and (iv) electromagnetic force $(F_{q})$. To ensure that the model remains in stable equilibrium, the combined effects of these forces must be balanced. In this study, to analyse the stability, we utilise the generalised form of the TOV equation [69, 70], as expressed below:
$-\frac{M_{G}(\rho+p_{r})}{r^{2}}e^{\lambda-\nu}-\frac{dp_{r}}{dr}+\frac{2\Delta}{r}+\frac{q}{4\pi r^{4}}\frac{dq}{dr}=0.$ (45)
The active gravitational mass, denoted by $M_{G}$ as given in Eq. (45) which can be derived using the mass formula as proposed by Tolman-Whittaker [82] in the following form:
@@ 6.1 TOV equation in a generalised form in presence of charge
$-\nu^{\prime}(\rho+p_{r})-\frac{dp_{r}}{dr}+\frac{2\Delta}{r}+\frac{q}{4\pi r^{4}}\frac{dq}{dr}=0.$ (47)
@@
Figure 12: Radial variation of different forces for different charge intensities $(q_{0})$. Here, Green, Red, Blue and Magenta lines represent $F_{g}$, $F_{h}$, $F_{a}$ and $F_{q}$ respectively. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@ 6.2 Herrera cracking condition
According to Herrera [83], the term ”Cracking” within a fluid sphere of a self-gravitating compact object represents the total radial forces originated, each with different signs, when perturbations are introduced into the system. If the EoS is taken into account, cracking can arise from two scenarios, namely, i) a local anisotropic fluid, or ii) a slowly contracting and radiating perfect fluid. Utilising this cracking condition, Abreu et al. [84] explored the impact of local anisotropy on the distribution of local anisotropic fluid. They demonstrated that when perturbations exist in the system, potential stability can be established through the difference between the squares of radial $(v_{r}^{2})$ and tangential $(v_{t}^{2})$ sound velocities expressed as:
$0\leq|v_{t}^{2}-v_{r}^{2}|\leq 1$ (52)
A stellar model that adheres to this relation is referred to as a stable structure. From Fig. (13), we note that the Abreu inequality is well maintained throughout the interior of the stellar configuration within the parameter space used here.
Figure 13: Radial variation of $|v_{r}^{2}-v_{t}^{2}|$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines indicate that $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@ 6.3 Adiabatic index
The adiabatic index $(\Gamma)$ signifies the rigidity of the EoS at a particular density. In the context of a relativistic anisotropic stellar structure, the adiabatic index is represented as:
@@
$\Gamma>\Gamma^{\prime},$ (54)
@@
$\Gamma^{\prime}=\frac{4}{3}-\Bigg{[}\frac{4}{3}\frac{(p_{r}-p_{t})}{|p^{\prime}_{r}|r}\Bigg{]}_{max}.$ (55)
Now, to determine the anisotropic limit, we have evaluated $\Gamma^{\prime}$ for different charge intensities $(q_{0})$ and the composite results are plotted in Fig. 14.
Figure 14: Adiabatic index $(\Gamma)$ vs r plot for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines indicate $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
In our model, it is evident from Fig. 14 that the value of $\Gamma$ consistently exceeds the threshold limit $\Gamma^{\prime}$ and hence the present model is stable under the consideration of adiabatic index variation.
@@ 6.4 Zel’dovich condition
Figure 15: Radial variation of $(\Omega_{r})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
Figure 16: Radial variation of adiabatic index $(\Omega_{t})$ for different charge intensities $(q_{0})$. The solid, dashed and dot-dashed lines represent $q_{0}=0.0002,~{}0.0004$ and $0.0006~{}Km^{-2}$ respectively.
@@ 7 Discussion
Finally, we summarise the major findings of the present paper and our efforts to mimic a charged compact stellar configuration in the context of symmetric teleparallel $f(Q)$ gravity. Taking into account a spherically symmetric space-time, we combine the KB metric ansatz [62], with a linear form of the $f(Q)$ action, which allows us to formulate a singularity-free solution of the EFE within the framework of $f(Q)$ theory. This solution incorporates a specific form of charge, denoted as $q(r)=q_{0}r^{3}$, where $q_{0}=0$ illustrates an uncharged stellar configuration.
Within this theoretical construct, we have selected PSR J0740+6620 [73] as a potential candidate to study its properties in presence of charge. We analyse and represent our results, both numerically and graphically, by varying the charge intensity parameter $(q_{0})$. It has been shown in Refs. [63, 64] that the modified theories of gravity lead to alterations in the boundary conditions of GR. This paper addresses these changes by assessing the continuity of both the induced metric as well as the extrinsic curvature tensors at the boundary of the stellar structure. We have used, (i) the matching condition of interior solutions with that of Reissner-Nordström [65, 66] exterior solution to compute the extrinsic curvature tensors, as obtained in Eqs. (39) and (40), and (ii) the condition of nullification of radial pressure at the stellar surface, to compute the necessary constant present in the model. We have noted the following key characteristic features:
@@
    Radial variation of the thermodynamic properties of a stellar structure, such as the energy density $(\rho)$, radial $(p_{r})$ and tangential $(p_{t})$ pressure profiles are represented in Figs. 1, 2, 3. As the charge intensity increases, the energy density and pressure profiles decrease. However, the monotonically decreasing nature of the characteristic profiles is well preserved in this model. Additionally, we may note that the increase in charge within a compact star generates electrostatic repulsion, which acts against the inward gravitational force, leading to a decrease in both energy density and pressure profiles. This effect occurs because the outward force from the accumulated charge reduces the compressive requirement to maintain high densities and pressures, particularly in the star’s core. As a result, the matter distribution shifts to balance the electric repulsion and gravitational forces, yielding lower density and pressure profiles, as illustrated in Figs. 1, 2, and 3. Furthermore, the radial variation of anisotropy is illustrated in Fig. 4. Fig. 4 indicates that as the charge intensity $(q_{0})$ increases, the electric field generates an additional outward force that aids in balancing the pressure distribution. This outward force partially offsets the strong inward gravitational pull, thereby reducing pressure anisotropy. Moreover, from Fig. 4, we note that, $\Delta>0$, which indicates a repulsive anisotropic behaviour. This nature further indicates a more massive stellar structure. We also note that at the centre $(r\rightarrow 0)$, $\Delta\rightarrow 0$, i.e., $p_{r}=p_{t}$, which is another viable condition in favour of stable compact star modeling. The radial variations of pressure and density gradients are negative, as shown in Figs. 5, 6 and 7, which aids in the maximisation of the core density. Hence, it may be concluded that the physical parameters are consistent throughout the stellar structure.
@@
    Obtaining the maximum mass through TOV equations depends on the particular choice of the EoS. In the present scenario, we have obtained the suitable EoS by maximising the radial sound velocity at the stellar centre, i.e., $\frac{dp_{r}}{d\rho}|_{r=0}=1$. Moreover, we have obtained the best fit EoS in this model, which are tabulated in Table 1 for different values of charge intensity $(q_{0})$. Using the EoS as tabulated in Table 1, we have numerically evaluated the maximum mass by solving the TOV equations, and Fig. 8 depicts the mass-radius plot. The maximum masses and radii for different charge intensities are listed in Table 2. Solution of TOV equation in this present model predicts that the maximum mass attainable in this context is $2.92~{}M_{\odot}$ for charge intensity $q_{0}=0.0009~{}Km^{-2}$. Moreover, in this model the radii of different pulsars and lighter component of GW190814 event have been predicted and are tabulated in Table 3. It is noted that our model complies with the observational results as well as predictions from other theoretical models [91, 92, 93, 94, 95, 96]. As for example, the predicted radius of lighter component of GW190814 event from our model is 11.82 Km which is very close to the value of $11.76^{+0.14}_{-0.19}$ Km predicted by Maurya et al. [96] using the Einstein-Gauss-Bonnet theory of gravity. It is noted that the model is valid up to the value of charge intensity $q_{0}\leq 0.0009~{}Km^{-2}$. Beyond this value, the model does not permit physically viable results.
@@
    Figs. 9 and 10 show the variations of radial $(v_{r}^{2})$ and tangential $(v_{t}^{2})$ sound velocities with the radial distance.Notably, with increasing charge intensity, both $v_{r}^{2}$ and $v_{t}^{2}$ decrease within the stellar boundary. However, it is evident that both the velocities are well-regulated, as they do not surpass the speed of light and comply with the necessary causality condition within the constrained value of parameters.
@@
    Considering the presence of charge, we have studied the necessary and sufficient energy conditions for a well-behaved compact star model within the framework of $f(Q)$ gravity, which is demonstrated in Fig. 11. We note that, all the energy conditions are fulfilled within the stellar structure for the parameter space used here.
@@ 8 Acknowledgments
DB is thankful to the Department of Science and Technology (DST), Govt. of India, for providing the fellowship vide no: DST/INSPIRE Fellowship/2021/IF210761. PKC gratefully acknowledges support from IUCAA, Pune, India under Visiting Associateship programme.